OpenStudy (goformit100):

The area of a rectangular piece of cardboard is 36 sq cm and its length is 9 cm. What is the width of the cardboard?

OpenStudy (anonymous):

4

OpenStudy (whpalmer4):

\[A = L*W\]\[W = A/L\]